OpenStudy (anonymous):

Need help with 3 math questions

OpenStudy (anonymous):

will fan and metal

OpenStudy (igreen):

^

OpenStudy (igreen):

\(\sf \dfrac{1}{8} a - \dfrac{3}{4} = \dfrac{1}{2}\) We want to keep the variable on one side of the equation, which means eliminating the \(\sf \dfrac{3}{4}\). To do that we have to perform the opposite operation.

OpenStudy (igreen):

Which, in this case is adding, since it's being subtracted. Make sense?

OpenStudy (anonymous):

is that the answer

OpenStudy (anonymous):

whats the answer

OpenStudy (anonymous):

Hello?? I need help

OpenStudy (anonymous):

@AihberKhan